OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 a. 2 b. 3 c. no solution d. all real numbers

OpenStudy (anonymous):

plz explain

OpenStudy (mertsj):

1. Remove the parentheses first using the distributive property.

OpenStudy (anonymous):

ok -2x - 6 = -2x - 2 -4

OpenStudy (mertsj):

Good job. Now collect the like terms on the right.

OpenStudy (anonymous):

ok is it -2x - 2 in the right side

OpenStudy (mertsj):

What is -2-4?

OpenStudy (anonymous):

oh, yeah -6

OpenStudy (mertsj):

So what is your equation now?

OpenStudy (anonymous):

-2x - 6 = -2x - 6

OpenStudy (mertsj):

Oh my land!!! Look at that!! A number is equal to itself. I think that is true for all numbers, don't you?

OpenStudy (anonymous):

oh, thx missed that thx alot

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

wait @Mertsj do you know what no solution means? can you give me an example

OpenStudy (anonymous):

4.1x - 6.2 = 2

OpenStudy (mertsj):

If you had an equation such as this? 2x+3=2x-8

OpenStudy (anonymous):

So... Your answer?

OpenStudy (mertsj):

Then you would subtract 2x from both sides and you would get 3=-8 which can never be true. So there is no value of x that will make that sentence true and hence there is no solution.

OpenStudy (anonymous):

thxs alot, i have troble sometimes firguring out my 5th math

OpenStudy (mertsj):

Keep working sweetheart. You'll get it.

OpenStudy (anonymous):

yup thxs

OpenStudy (mertsj):

yw